# Data-Driven Dynamic Factor Modeling via Manifold Learning

### v1

@@ 3.2 Data-Driven Embedding
Following [singer2008nonlinear], we describe the anisotropic diffusion maps algorithm which is used to uncover the evolution of the latent coordinates. Let $Z=H(\theta)$ be a $C^{2}\cap\mathbb{L}^{2}(\mu)$ mapping of $\theta$, and suppose we are given a data set of $N+1$ discrete-time observations of $Z$, denoted $(z(t_{i}))_{i=0}^{N}$. Depending on the application at hand, and the size and quality of the data sets, we may for instance take $Z=X$, $Z=Y$, or $Z=(X,Y)$. To fix notation, we consider the situation here where $Z=(X,Y)$ so that $H=(F,G):\mathbb{R}^{d}\to\mathbb{R}^{m+n}$. We assume that the data has been mean-centered.
Motivated by the Taylor approximation (3.7) in the previous subsection, we consider the modified Mahalanobis distance between $z_{i}$ and $z_{j}$:
@@
where $C(t_{i})$ denotes the instantaneous quadratic covariation matrix of $z(t_{i})$. In Section 5, we approximate $C(t_{i})$ by its sample average in a trailing window. For strictly positive financial data, one can also use an estimator such as the one presented in [chesney_estimating_1995].
We construct a distance matrix for the data $\mathbf{D}\in\mathbb{R}^{N+1}\times\mathbb{R}^{N+1}$ with entries $\mathbf{D}_{ij}=d(z(t_{i}),z(t_{j}))$. We pass the entries of $\mathbf{D}$ through the Radial Basis Function (RBF) kernel $\exp\left(-d(\cdot,\cdot)/2\varepsilon\right)$ with parameter $\varepsilon>0$ to obtain an adjacency matrix $\mathbf{W}$:
@@
A popular choice for practitioners is to choose $\varepsilon$ as the median of all distances between data points (see [shnitzer2020diffusion] or Subsection 6.2). Another methodology (see [singer_detecting_2009, Figure 3]) is to choose $\varepsilon$ in a region where the Log-Log plot of the sum of all entries of $\mathbf{W}$ versus $\varepsilon$ is linear. In any case, $\mathbf{W}$ is then normalized to create a stochastic matrix $\mathbf{P}$ (which is in fact doubly-stochastic as it is symmetric):
@@
In Theorem 4.3, we show that the graph Laplacian $\mathbf{P}-I$ approximates the Fokker–Planck operator $\mathscr{L}$ in the following sense: if $(z(t_{i}))_{i=0}^{N}=(H(\theta(t_{i})))_{i=0}^{N}$ where $\theta$ satisfies the dynamics in Assumption 1, and if $f:\mathbb{R}^{m+n}\to\mathbb{R}$ is such that $f\circ H\in\mathbb{L}^{2}(\mu)$ then the $i$th row of $(\mathbf{P}-I)$ applied to the vector $[f(z(t_{0})),\dots,f(z(t_{N}))]$ satisfies
@@
As our data is mean-centered, we discard the first eigenvector (a constant vector corresponding to eigenvalue $1$) and keep eigenvectors 1 through $\ell$, for some $\ell>0$. Practitioners commonly choose $\ell$ to correspond to the eigenvalue preceding the largest drop-off $\kappa_{i}-\kappa_{i+1}$ (see, for instance, Subsection 6.2 or [coifman2008diffusion, Subsection 3.2]). In the context of Example 3.3, a higher value of $\ell$ corresponds to a set of polynomials with higher maximum degree.
@@
Following, [shnitzer2020diffusion], we approximate the observed data by a lifting operator $z(t_{i})=\mathbf{H}\psi(t_{i})$, which is given in coordinates by
@@
Notice that $\langle\cdot,\cdot\rangle_{\text{data}}$ is defined in [shnitzer2020diffusion] without the normalization by $N+1$. Note also that the diffusion coordinates themselves are of order $O(N^{-1/2})$, as they are the coordinates of $\mathbb{R}^{N+1}$-orthonormal vectors. These scaling factors are of no consequence for the state-space model presented in the next subsection, since the diffusion coordinates are treated as hidden variables and may be rescaled as one wishes.
@@
Before turning to our data-driven methodology, we return to Example 3.3 and illustrate the use of diffusion maps on synthetic data. We perform a simulation of the observation vector $z(t)=(x(t),y(t))$ composed of 500 discrete-time samples of the processes $X$ and $Y$, respectively. The data has been mean-centered and the covariation process is approximated using a windowed covariance matrix, with a trailing window of size 50. Since we have access to the true $\theta$ and the ground-truth Hermite polynomial eigenfunctions, we can assess how well the diffusion coordinates recover the eigenfunctions of $\mathscr{L}$, as well as the reconstruction quality of the $X$ and $Y$ processes via the lifting operator $\mathbf{H}$, using the dominant $10$ diffusion map eigenvectors. Figures 2a and 2b below display the reconstructed $X$ and $Y$ processes both in terms of time series and hypersurfaces.
(a) True $X_{t}$ (left) and $Y_{t}$ (right) surfaces as a function of $\theta_{t}$ and stem plots of the lifted reconstructions via $\mathbf{H}\psi(t)$.
(b) True $X_{t}$ (left) and $Y_{t}$ (right) as a function of time and the lifted time series via $\mathbf{H}\psi(t)$.
Similarly, Figures 2c and 2d below display two ground-truth Hermite eigenfunctions of the backward Fokker-Planck operator $\mathscr{L}$ and their approximation via the diffusion map eigenvectors, both in terms of time series and hypersurfaces. The observed fit is satisfactory in these cases; however, we note that the diffusion maps algorithm may be learning a different basis than these Hermite eigenfunctions (even restricting to polynomials, there are an uncountable number of such bases due to the rotational invariance of the Gaussian density).
(c) Hermite polynomials $h_{1,0}$ (left) and $h_{1,3}$ (right) surfaces as a function of $\theta_{t}$ alongside stem plots of $\psi_{1}(t)$ and $\psi_{3}(t)$ respectively.
(d) Hermite polynomials $h_{1,0}$ (left) and $h_{1,3}$ (right) as a function of time against the time series $\psi_{1}(t)$ and $\psi_{3}(t)$ respectively.
@@
3.3 Joint Diffusion Kalman Filter

### v2

@@ 3.2 Data-Driven Embedding
We outline here the anisotropic diffusion maps algorithm for learning the eigenvalues and eigenfunctions of the generator $\mathscr{L}$ from discrete observations. The algorithm constructs a graph Laplacian from data that approximates $I+\varepsilon\mathscr{L}$ for small $\varepsilon>0$, enabling recovery of the generator’s spectral properties without parametric assumptions.
Let $Z=H(\theta)$ be a $C^{2}\cap\mathbb{L}^{2}(\mu)$ mapping of $\theta$, and suppose we are given a data set of $N+1$ discrete-time observations of $Z$, denoted $(z(t_{i}))_{i=0}^{N}$. Depending on the application at hand, and the size and quality of the data sets, we may for instance take $Z=X$, $Z=Y$, or $Z=(X,Y)$. To fix notation, we consider the situation here where $Z=(X,Y)$ so that $H=(F,G):\mathbb{R}^{d}\to\mathbb{R}^{m+n}$. We assume that the data has been mean-centered.
Although we do not observe $\theta$ directly, distances in the latent space can be approximated using observations of $Z$. Motivated by the Taylor approximation (3.7) in the previous subsection, we consider the squared modified Mahalanobis distance between $z_{i}$ and $z_{j}$:
@@
where $C(t_{i})$ denotes the instantaneous quadratic covariation matrix of $z(t_{i})$. This approximates $\|\theta(t_{i})-\theta(t_{j})\|^{2}$ up to $O(\|z(t_{i})-z(t_{j})\|^{4})$ by equation (3.7). In Section 6, we approximate $C(t_{i})$ by its sample average in a trailing window. For strictly positive financial data, one can also use an estimator such as the one presented in [12].
We construct a squared distance matrix for the data $\mathbf{D}\in\mathbb{R}^{N+1}\times\mathbb{R}^{N+1}$ with entries
$\mathbf{D}_{ij}=d(z(t_{i}),z(t_{j}))$}. We pass the entries of $\mathbf{D}$ through the Radial Basis Function (RBF) kernel $\exp\left(-d(\cdot,\cdot)/2\varepsilon\right)$ with parameter $\varepsilon>0$ to obtain an adjacency matrix $\mathbf{W}$:
@@
A popular choice for practitioners is to choose $\varepsilon$ as the median of all distances between data points (see [36] or Subsection 7.2). Another methodology (see [40, Figure 3 ]) is to choose $\varepsilon$ in a region where the Log-Log plot of the sum of all entries of $\mathbf{W}$ versus $\varepsilon$ is linear. In any case, $\mathbf{W}$ is then normalized to create a stochastic matrix $\mathbf{P}$ (which is in fact doubly-stochastic as it is symmetric):
@@
In Theorem 4.3, we show that the graph Laplacian $\mathbf{P}-I$ approximates the Fokker–Planck operator $\mathscr{L}$ in the following sense: if $(z(t_{i}))_{i=0}^{N}=(H(\theta(t_{i})))_{i=0}^{N}$ where $\theta$ satisfies the dynamics in Assumption 1, and if $f:\mathbb{R}^{m+n}\to\mathbb{R}$ is such that $f\circ H\in\mathbb{L}^{2}(\mu)$ then the $i$th row of $(\mathbf{P}-I)$ applied to the vector $f(z(t_{0})),\dots,f(z(t_{N}))]$ satisfies
@@
As our data is mean-centered, we discard the first eigenvector (a constant vector corresponding to eigenvalue $1$) and keep eigenvectors 1 through $\ell$, for some $\ell>0$. Practitioners commonly choose $\ell$ to correspond to the eigenvalue preceding the largest drop-off $\kappa_{i}-\kappa_{i+1}$ (see, for instance, Subsection 7.2 or [13, Subsection 3.2 ]). In the context of Example 3.3, a higher value of $\ell$ corresponds to a set of polynomials with higher maximum degree.
@@
Following, [36], we approximate the observed data by a lifting operator $z(t_{i})=\mathbf{H}\psi(t_{i})$, which is given in coordinates by
@@
Notice that $\langle\cdot,\cdot\rangle_{\text{data}}$ is defined in [36] without the normalization by $N+1$. Note also that the diffusion coordinates themselves are of order $O(N^{-1/2})$, as they are the coordinates of $\mathbb{R}^{N+1}$-orthonormal vectors. These scaling factors are of no consequence for the state-space model presented in Section 5, since the diffusion coordinates are treated as hidden variables and may be rescaled as one wishes.
@@
Before turning to our theoretical results and data-driven methodology, we return to Example 3.3 and illustrate the use of diffusion maps on synthetic data. We perform a simulation of the observation vector $z(t)=(x(t),y(t))$ composed of 500 discrete-time samples of the processes $X$ and $Y$, respectively. The data has been mean-centered and the covariation process is approximated using a windowed covariance matrix, with a trailing window of size 50. Since we have access to the true $\theta$ and the ground-truth Hermite polynomial eigenfunctions, we can assess how well the diffusion coordinates recover the eigenfunctions of $\mathscr{L}$, as well as the reconstruction quality of the $X$ and $Y$ processes via the lifting operator $\mathbf{H}$, using the dominant $10$ diffusion map eigenvectors. Figures 1a and 1b display the reconstructed $X$ and $Y$ processes both in terms of time series and hypersurfaces.
(a) True surfaces $X_{t}=(\theta_{t}^{1})^{2}+(\theta_{t}^{2})^{2}$ and $Y_{t}=\theta_{t}^{1}$ (colored by value) overlaid with lifted reconstructions $\mathbf{H}\psi(t)$ (stem plots).
(b) True $X_{t}$ (left) and $Y_{t}$ (right) are plotted as functions of time alongside the lifted time series via $\mathbf{H}\psi(t)$.
Figure 1: Comparisons showing that $\ell=10$ diffusion coordinates accurately reconstruct the OU and CIR dynamics from Example $3.$
Similarly, Figures 2a and 2b below display two ground-truth Hermite eigenfunctions of the backward Fokker-Planck operator $\mathscr{L}$ and their approximation via the diffusion map eigenvectors, both in terms of time series and hypersurfaces. The observed fit is satisfactory in these cases; however, we note that the diffusion maps algorithm may be learning a different basis than these Hermite eigenfunctions (even restricting to polynomials, there are an uncountable number of such bases due to the rotational invariance of the Gaussian density).
(a) Hermite polynomials $h_{1,0}$ (left) and $h_{1,3}$ (right) surfaces as a function of $\theta_{t}$ alongside stem plots of $\psi_{1}(t)$ and $\psi_{3}(t)$ respectively.
(b) Hermite polynomials $h_{1,0}$ (left) and $h_{1,3}$ (right) as a function of time against the time series $\psi_{1}(t)$ and $\psi_{3}(t)$ respectively.
@@
4 Convergence of Averages and Robustness of Approximations
Having described the diffusion maps algorithm and illustrated its performance on synthetic data, in this section we consider the theoretical foundations. We establish three key results: (i) the graph Laplacian constructed from time series data converges to the true generator (Theorem 4.3), (ii) the lifting operator converges to the $\mathbb{L}^{2}(\mu)$ projection (Proposition 4.2), and (iii) the linear approximation of eigenfunction dynamics is robust under ergodic averaging (Section 4.3). These results rigorously justify the Kalman filtering framework we develop in Section 5.
The works by [36, 37] leverage the algorithm put forth in [39] to learn the eigenvalues of $\mathscr{L}$, to estimate the dynamics of the eigenfunctions, and to approximate the observations using a linear lift. The Kalman filtering framework of [36] is the motivation for the Joint Diffusion Kalman Filter and conditional sampling procedure we subsequently develop in Section 5. However, there are some mathematical questions that first need to be addressed in order to rigorously justify this procedure. Under the assumptions in Subsection 3.1, we show in this section that it is admissible to learn the eigenvalues and the lift from time series data. In particular, a Poincaré Inequality (Assumption 3) implies that initialization errors are forgotten, and time averages converge (see Subsection 4.1). In Subsection 4.2, we bound the approximation error of the eigenfunctions of $\mathscr{L}$ by the diffusion map eigenvectors using a concentration inequality. Furthermore, in Subsection 4.3 we quantify the error of approximating the eigenfunction dynamics using linear diffusions, demonstrating robustness.
$S_{N}=\frac{1}{N+1}\sum_{i=0}^{N}V(\theta(t_{i})).$ (4.1)
For instance, in the setting of [38], $(\theta(t_{i}))_{i=0}^{N}$ are independent uniform samples from a compact Riemannian manifold $\mathcal{M}$, and the graph Laplacian is shown to approximate the Laplace–Beltrami operator $\Delta_{\mathcal{M}}$, that is, the following holds for smooth $f$ and a point $\theta(t_{i})$ fixed:
In our setting, the summands in (4.1) arise not from independent samples, but from a sampled Langevin diffusion, which yields a Markov chain. In this case, the convergence of $S_{N}$ to the expectation $\mathbb{E}_{\mu}[V]$ as $N\to\infty$ is an ergodic property, which is guaranteed for $V\in\mathbb{L}^{2}(\mu)$ under our assumption of a Poincaré inequality or (equivalently) a spectral gap. If the Langevin diffusion is started in equilibrium, i.e. $\theta_{0}\sim\mu$, then $(\theta(t_{i}))_{i=0}^{N}$ is a stationary sequence and $\mathbb{E}_{\theta_{0}\sim\mu}[S_{N}]=\mathbb{E}_{\mu}[V]$ for any $N\geq 0$. In this case, the fluctuations $\sqrt{N}S_{N}$ converge as $N\to\infty$ to a Gaussian distribution by the celebrated Central Limit Theorem of [27]. A fast rate of convergence to equilibrium for Markov chains is a desirable property in the context of Markov chain Monte Carlo algorithms, and many central limit theorem results are collected in the survey paper [34]. For continuous time averages, we refer the reader to [9], where the trend to equilibrium is studied under weaker assumptions than ours (for instance, some laws with sub-exponential tails are shown to be ergodic). In general, the rate of convergence depends both on the integrability of $V$ and the weight of the tails of $\mu$. To fix notation, throughout the remainder of this section we consider a constant fixed step size $\Delta t=t_{i}-t_{i-1}$ and set $t_{0}=0$.
4.1 Short-Term Equilibration and Long-Term Convergence
where we have used Assumption 3 for the inequality. The differential form of Grönwall’s Inequality yields $\|P_{t}f\|_{\mathbb{L}^{2}(\mu)}^{2}\leq\exp(-2t/C)\|f\|_{\mathbb{L}^{2}(\mu)}^{2}$. Hence, $P_{t}f$ converges to equilibrium in $\mathbb{L}^{2}$ exponentially fast. In fact, this property is equivalent to a Poincaré Inequality (see for instance [2, Theorem 4.2.5 ]). In the next lemma, we exploit this fact to bound the deviation of observables from their values at equilibrium when $\theta_{0}$ is initialized at some $\mu_{0}$ which is absolutely continuous with respect to the equilibrium measure $\mu$. In practice, this error can be diminished by discarding an initial burn-in period (as we have done in Subsection 6.2). Similar results hold in other $\mathbb{L}^{p}$-spaces by Riesz–Thorin Interpolation (see, for instance, [35, Proposition 3.17 ] and the application to MCMC methods in [19, Theorem 12 ]), but we restrict to the $\mathbb{L}^{2}$ case to reduce notation.
Lemma 4.1.
Proof.
Since $\mu_{t}$ is absolutely continuous with respect to $\mu$ (as the Langevin diffusion preserves absolute continuity), we can write
$\mathbb{E}_{\theta_{0}\sim\mu_{0}}\left[f(\theta_{t})\right]$ $=\int_{\mathbb{R}^{d}}f(w)\mu_{t}(\mathrm{d}w)=\int_{\mathbb{R}^{d}}f(w)\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\mu}(w)\mu(\mathrm{d}w)$
$=\int_{\mathbb{R}^{d}}f(w)\left(1+\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\mu}(w)-1\right)\mu(\mathrm{d}w)$
$=\mathbb{E}_{\mu}[f]+\int_{\mathbb{R}^{d}}f(w)\left(\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\mu}(w)-1\right)\mu(\mathrm{d}w).$
Applying Hölder’s inequality to the second term yields
$\left|\int_{\mathbb{R}^{d}}f(w)\left(\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\mu}(w)-1\right)\mu(\mathrm{d}w)\right|$ $\leq\|f\|_{\mathbb{L}^{2}(\mu)}\left\|\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\mu}-1\right\|_{\mathbb{L}^{2}(\mu)}.$
The density evolution $\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\mu}$ satisfies $\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\mu}=P_{t}\left(\frac{\mathrm{d}\mu_{0}}{\mathrm{d}\mu}\right)$ where $P_{t}$ is the Markov semigroup associated with $\theta$. Since $\mathbb{E}_{\mu}\left[\frac{\mathrm{d}\mu_{0}}{\mathrm{d}\mu}\right]=1$, we have $\mathbb{E}_{\mu}\left[\frac{\mathrm{d}\mu_{0}}{\mathrm{d}\mu}-1\right]=0$, and thus
$\left\|\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\mu}-1\right\|_{\mathbb{L}^{2}(\mu)}$ $=\left\|P_{t}\left(\frac{\mathrm{d}\mu_{0}}{\mathrm{d}\mu}-1\right)\right\|_{\mathbb{L}^{2}(\mu)}.$
By the $\mathbb{L}^{2}$-exponential convergence of $P_{t}$ established via Grönwall’s inequality in the text preceding this lemma (with rate $\lambda_{1}=1/C$ where $C$ is the Poincaré constant), we obtain
$\left\|P_{t}\left(\frac{\mathrm{d}\mu_{0}}{\mathrm{d}\mu}-1\right)\right\|_{\mathbb{L}^{2}(\mu)}$ $\leq\exp\left(-\lambda_{1}t\right)\left\|\frac{\mathrm{d}\mu_{0}}{\mathrm{d}\mu}-1\right\|_{\mathbb{L}^{2}(\mu)}.$
Combining these estimates yields the desired result. ∎
The first sum concerns the approximation of the eigenfunctions of $\mathscr{L}$ by diffusion coordinates, which is addressed in the next subsection. Focusing on the second term, we shall apply the Kipnis–Varadhan Central Limit Theorem [27]. We state the result with $\theta_{0}\sim\mu$; however, under suitable assumptions on $F^{j}\varphi_{k}$ similar results hold even when the Markov chain is started from a deterministic initial condition (see, for instance, [16, 17, 15, 9]).
Proposition 4.2 ([27]).
holds, where $\sigma^{2}$ satisfies
and $\rho_{F^{j}\varphi_{k}}$ is the spectral measure of $F^{j}\varphi_{k}$.
4.2 Error Analysis of the Graph Laplacian
The ADM approach of [39] relies on the approximation of the Fokker–Planck operator by the graph Laplacian. Suppose $(\theta(t_{i}))_{i=0}^{N}$ are i.i.d. samples from $\mu$, let $f\in\mathbb{L}^{2}(\mu)$, and set $z(t_{i})=H(\theta_{i})$. Then an analogous argument to (4.2) (see [39, Equation (30) ] along with extra accounting for a factor of 2 from differences in some definitions given there) yields
$\sum_{j=0}^{N}\mathbf{P}_{ij}f(z(t_{j}))\to\frac{\mathbb{E}_{\mu}\left[\exp\left(-\frac{\|\theta-\theta(t_{i})\|^{2}}{2\varepsilon}\right)f\circ H(\theta)\right]}{\mathbb{E}_{\mu}\left[\exp\left(-\frac{\|\theta-\theta(t_{i})\|^{2}}{2\varepsilon}\right)\right]}=f\circ H(\theta(t_{i}))+\varepsilon\mathscr{L}(f\circ H)(\theta(t_{i}))+O(\varepsilon^{2})$
Setting $w=\theta(t_{i})$, we have that
Theorem 4.3.
where $C$ depends on $\|u\|_{\infty}$.
Remark 4.4.
Proof.
We adapt the argument from [38, Section 3 ] to samples $(\theta(t_{i}))_{i=0}^{N}$ from the Langevin diffusion (3.1). In this case, $(\theta(t_{i}))_{i=0}^{N}$ forms a reversible Markov chain with spectral gap $1-\exp(-\lambda_{1}\Delta t)$. We will show an upper bound for the right tail in the desired inequality. The left tail follows similarly and the statement of the theorem holds by a union bound. If we let
then we see that
Note that $\mathbb{E}_{\mu}[V]=0$ and $V$ is bounded in absolute value by some constant $C$ which depends on $\|u\|_{\infty}$. We seek to apply a concentration inequality which will capture the second moment of $V$ in the variance proxy. Since $(\theta(t_{i}))_{i=0}^{N}$ forms a reversible Markov chain with respect to $\mu$ and spectral gap $1-\exp(-\lambda_{1}\Delta t)$ (where $\lambda_{1}$ is the spectral gap of -$\mathscr{L}$ given by the Poincaré inequality in Assumption 3), we can apply Bernstein’s inequality for Markov chains [32, Theorem 3.3]. This inequality requires bounding the second moment $\mathbb{E}_{\mu}[V^{2}]$, which serves as a variance proxy in the non-i.i.d. setting. The inequality gives
A lengthy computation (which is relegated to Appendix D) shows that
and the conclusion of the theorem follows. ∎
where $B^{i}$ is a Brownian motion, and for $i\neq j$ we have
Corollary 4.5.
Proof.
For $u\geq 0$, let $\mu_{u}$ denote the law of $\theta_{u}$. We have that
In [36], the gradients $\|\nabla\varphi_{i}(\theta_{t})\|_{2}$, for $i=1,\dots,\ell$ are assumed to be constant, so that the SDE for $\varphi_{i}(\theta_{t})$ is linear-Gaussian and a Kalman filtering framework may be applied. As another corollary of Lemma 4.1, we show next that the approximation error between the contractive dynamics (4.4) and a linear SDE can remain small. For $i=1,\dots,\ell$, let $\xi^{1},\dots,\xi^{\ell}$ satisfy
Corollary 4.6.
where
Remark 4.7.
Proof.
By Itô’s formula, the squared error at time $t\geq 0$ satisfies
For $u\geq 0$, let $\mu_{u}$ denote the law of $\theta_{u}$. By Lemma 4.1, we have
The final result follows from using the above inequality in (4.7) and applying the integral form of Grönwall’s Inequality. ∎
5 Joint Diffusion Kalman Filter
In Section 4, we have developed the main results to justify a state space model in terms of the diffusion coordinate embeddings: 1) the generator eigenfunction dynamics can be approximated by a linear SDE with drift equal to the corresponding eigenvalue (Subsection 4.3); 2) the diffusion coordinates $\psi$ approximate the eigenfunctions of the generator associated with the latent state $\theta$ (Subsection 4.2); 3) the lifting operator learned from time series data converges to the true $\mathbb{L}^{2}(\mu)$ projection of the observation map onto the generator eigenfunctions (Subsection 4.1). In this section, we describe our linear state space approximation for the dynamics of the diffusion coordinates, leveraging these theoretical results developed in Section 4.